# Hey Dashboard!: Supporting Voice, Text, and Pointing Modalities in Dashboard Onboarding

### v1

@@ Abstract
Visualization dashboards are regularly used for data exploration and analysis, but their complex interactions and interlinked views often require time-consuming onboarding sessions from dashboard authors. Preparing these onboarding materials is labor-intensive and requires manual updates when dashboards change. Recent advances in multimodal interaction powered by large language models (LLMs) provide ways to support self-guided onboarding. We present Diana (Dashboard Interactive Assistant for Navigation and Analysis), a multimodal dashboard assistant that helps users for navigation and guided analysis through chat, audio, and mouse-based interactions. Users can choose any interaction modality or a combination of them to onboard themselves on the dashboard. Each modality highlights relevant dashboard features to support user orientation. Unlike typical LLM systems that rely solely on text-based chat, Diana combines multiple modalities to provide explanations directly in the dashboard interface. We conducted a qualitative user study to understand the use of different modalities for different types of onboarding tasks and their complexities.
@@
Figure 1: Multimodal interaction for dashboard onboarding using LLMs. Using the process model for dashboard onboarding proposed by Dhanoa et al. [12], we extend it with multimodal interactions and features, powered by LLMs for onboarding a user.
1 Introduction
@@
Visualization dashboards have become essential tools for data exploration and decision-making across organizations [33]. Modern dashboards integrate multiple linked views, complex filtering mechanisms, and sophisticated interaction techniques to support analytical workflows [4]. However, this complexity creates a significant barrier: users often struggle to understand how to navigate and effectively use these dashboards without guidance [44, 7, 10]. Dashboard authors accordingly invest substantial time creating onboarding materials and conducting training sessions [12, 40], yet these resources quickly become outdated as dashboards evolve [13]. Existing automated onboarding solutions often fail to match the effectiveness of human experts who can provide contextual explanations, adapt to user questions, and demonstrate features using the actual data and visualizations present in the dashboard.
@@
We present Diana (Dashboard Interactive Assistant for Navigation and Analysis), a multimodal dashboard assistant that replicates the guidance provided by human onboarding experts. Unlike text-only chatbots or onboarding tutorials (static or dynamic [13]), Diana combines multiple interaction modalities—voice input and output, mouse, and keyboard-based interaction—to provide situated explanations [9] directly within the dashboard interface. The system employs a contextual radial menu that organizes dashboard features into logical structures, supports pointing gestures and lasso selection for highlighting specific charts or data points, includes a text-based natural language interface, and integrates speech recognition and synthesis for natural conversation. Diana delivers contextual help using the actual tasks, visualizations, and datasets from each specific dashboard, enabling users to learn through direct manipulation and exploration of real content.
@@
We implemented Diana as a web-based system using a GPT-powered agent and integrated with Microsoft Power BI. The system supports any Power BI dashboard with appropriate metadata annotations. It uses OpenAI’s Realtime API [2] for speech processing and provides real-time visual feedback through dashboard highlighting and overlays. Using the prototype, we conducted an in-depth qualitative user study with novices and expert dashboard users across hour-long sessions involving detailed tasks on varied, unfamiliar datasets and dashboards. The study design incorporated insights from three pilot studies to refine the experimental protocol. Results demonstrate that participants rapidly grasped Diana’s capabilities and consistently preferred multimodal interactions over single-modality approaches for onboarding tasks. Visual highlighting combined with voice-based questions and responses emerged as the dominant interaction pattern. All experts expressed strong interest in using Diana for their professional dashboard work.
@@
The contributions of this work are threefold: (1) the concept of a multimodal dashboard onboarding assistant that combines natural language, visual highlighting, and gestural interaction to replicate human-expert guidance; (2) the implementation of Diana, demonstrating how large language models can be integrated with dashboard environments to provide contextual, situated assistance; and (3) empirical findings from a qualitative user study revealing how novice and expert users employ different modalities for various onboarding tasks and their preferences for multimodal over unimodal interaction patterns.
2 Related Work
@@ 2.1 Visualization Dashboards
Visualization dashboards have become ubiquitous tools for data exploration and decision-making across diverse domains, yet their complexity often creates significant barriers for users. Few’s book on visualization dashboards [18] outlined the many fine points and varied purposes of dashboard design in industry. Sarikaya et al. [34] provided a comprehensive survey of what the scientific visualization community means by “dashboards,” revealing their diverse applications and the varied expectations users bring to these systems. Their work highlighted the fundamental challenge: dashboards must balance powerful functionality with usability.
@@
Multiple studies have documented the struggles users face when working with dashboard systems. Tory et al. [43] examined how users attempt to find their “data voice” when working with dashboards, documenting persistent difficulties in navigation, interpretation, and effective use of interactive features. Their findings reveal that even well-designed dashboards can overwhelm users who lack sufficient training. Similarly, Walchshofer et al. [44] observed the socio-technical challenges that arise when organizations transition to new dashboarding systems, noting that successful adoption depends heavily on both technical design and organizational support strategies.
@@
The visualization community has responded to these challenges by developing design frameworks and best practices. Bach et al. [4] identified recurring dashboard design patterns, providing a vocabulary for discussing effective dashboard layouts and interaction paradigms. Lin et al. [23] developed DMiner, which provides design recommendations for dashboards based on analysis of a large dataset of existing dashboards. Their system helps authors create more effective dashboards by suggesting layouts, chart types, and design patterns that have proven successful in similar contexts, representing a shift toward data-driven dashboard design that could reduce user confusion from the outset.
@@
Despite these advances in design guidance, the challenge remains: complex dashboards with multiple coordinated views and rich interactions require significant user investment to master. This persistent gap between dashboard capabilities and user proficiency motivates the need for more sophisticated onboarding and assistance mechanisms that can help users navigate these powerful but complex systems.
2.2 Onboarding for Dashboards
@@
Given the popularity of visualization dashboards, visualization onboarding has recently emerged as a research area, though solutions remain fragmented across different contexts and applications. Stoiber et al. [40] first characterized the onboarding space for individual visualizations, establishing foundational concepts and identifying the need for systematic approaches to help users learn visualization systems. Their work highlighted various onboarding methods ranging from simple tooltips to more sophisticated guided experiences.
@@
Building on these foundational concepts, Dhanoa et al. [12] proposed a systematic process model for dashboard onboarding, identifying key stages that users go through when learning new dashboard systems. This work provided a framework for understanding onboarding as a structured process rather than ad-hoc training.
@@
Recent advances have focused on automating onboarding creation using tours. Dhanoa et al. [13] introduced D-Tour, a system for semi-automatically generating interactive guided tours for visualization dashboards. Their approach reduces the manual effort required from dashboard authors while maintaining the ability to create engaging onboarding experiences. Similarly, Hoque and Sultanum [19] developed DashGuide, which enables the creation of interactive dashboard tours with minimal author input while allowing for manual refinement when needed.
@@
The comparative evaluation of different onboarding approaches has revealed important insights about their effectiveness. Stoiber et al. [41] systematically compared various onboarding methods, finding that interactive approaches generally outperform static documentation, though the optimal approach depends on user characteristics and dashboard complexity.
@@
Despite these advances, most existing onboarding solutions remain primarily visual and static, relying on annotations [15], tooltips, or guided tours that users follow passively. Chundury et al. [9] represent a notable exception with their contextual, in-situ help features for visual data interfaces, though their approach focuses on help rather than comprehensive onboarding. This gap between the interactive nature of modern dashboards and the static nature of most onboarding solutions motivates the need for more dynamic, responsive onboarding systems that can adapt to user behavior and provide assistance through multiple interaction modalities. In particular, current automated onboarding methods are still far inferior to the gold standard of onboarding: using a human instructor .
2.3 Multimodal Interaction with Data Visualizations
@@
The visualization community has long recognized the limitations of traditional mouse and keyboard interaction for complex data exploration tasks. Elmqvist et al. [16] articulated the concept of fluid interaction for information visualization, arguing that effective visualization systems should support seamless transitions between different interaction modes and reduce cognitive barriers between user intent and system response. This foundational work established fluidity as a key principle and identified the utility of multiple modalities for designing more natural visualization interfaces.
@@
Lee et al. [22] provided a comprehensive call to action, encouraging the visualization community to expand beyond mouse and keyboard to support more natural interactions. Their survey of alternative interaction modalities highlighted the potential for touch, speech, gesture, and other modalities to make visualization more accessible and expressive. This work catalyzed subsequent research into multimodal visualization interfaces.
@@
Early explorations of touch-based visualization interaction demonstrated the potential of direct manipulation using modalities beyond the mouse and keyboard. Baur et al. [6] introduced TouchWave, showing how kinetic multi-touch gestures could enable fluid manipulation of hierarchical stacked graphs. Their work revealed how touch affordances could support more intuitive interactions with complex visualization structures. Building on these foundations, Nielsen et al. [29] studied the use of touch and pen scribbling to support multidimensional visualization queries. Badam et al. [5] analyzed the affordances of different input modalities—touch, speech, proxemics, gestures, gaze, and wearables—for visual data exploration, providing guidelines for combining modalities to leverage their complementary strengths. Thompson et al. [42] developed Tangraphe, which employed single-hand, multi-touch gestures for network visualization exploration, demonstrating how naturalistic gestures could replace conventional pointer-based interactions.
@@
The integration of speech with visualization interaction has proven particularly promising for complex analytical tasks. Srinivasan and Stasko [37] examined what users naturally want to express through natural language when analyzing data, providing insights into the linguistic patterns that visualization systems should support. They later developed Orko [38], one of the first systems to combine voice and direct manipulation for network analysis, showing how speech could complement visual interaction for complex exploration tasks. These systems demonstrated that users could express analytical intents through conversational interfaces, though challenges remained in handling ambiguous queries and maintaining context across extended analysis sessions. Building on these foundations, Srinivasan and Setlur [36] developed BOLT, a natural language interface specifically designed for dashboard authoring, showing how conversational interaction could extend beyond data exploration to visualization creation itself.
@@
Empirical studies have provided crucial insights into user preferences and performance with different modality combinations. Saktheeswaran et al. [32] found that participants preferred combining multimodal over unimodal input when exploring network visualizations on large vertical displays, with touch and speech proving to be a particularly effective combination. Their findings challenged assumptions that users would prefer single-modality interaction for simplicity. In large display contexts, León et al. [28] found that participants preferred speech commands for 10 of 15 exploration tasks over touch, pen, and mid-air gestures, suggesting that speech excels particularly in scenarios where physical interaction is challenging.
@@
Recent work has focused on designing consistent multimodal experiences across different devices. Jakobsen et al. [20] explored how user position and movement relative to large displays could inform interaction design, showing how proxemic cues could enhance information visualization experiences through implicit spatial input. Srinivasan et al. [39] developed InChorus, which provides consistent multimodal interactions across tablet devices, demonstrating how design principles can maintain usability while adapting to different form factors.
2.4 Automatic Generation of Dashboards

### v2

# “Hey Dashboard!”: Supporting Voice, Text, and Pointing Modalities in Dashboard Onboarding using Large Language Models

@@ Abstract.
Visualization dashboards are regularly used for data exploration and analysis, but their complex interactions and interlinked views often require time-consuming onboarding sessions from dashboard authors. Preparing these onboarding materials is labor-intensive and requires manual updates when dashboards change. Recent advances in multimodal interaction powered by large language models (LLMs) provide ways to support self-guided onboarding. We present Diana (Dashboard Interactive Assistant for Navigation and Analysis), a multimodal dashboard assistant that helps users for navigation and guided analysis through chat, audio, and mouse-based interactions. Users can choose any interaction modality or a combination of them to onboard themselves on the dashboard. Each modality highlights relevant dashboard features to support user orientation. Unlike typical LLM systems that rely solely on text-based chat, Diana combines multiple modalities to provide explanations directly in the dashboard interface. We conducted a comparative qualitative user study to understand the use of different modalities for different types of onboarding tasks and their complexities.
visualization, dashboards, multimodal interactions, large language models, onboarding, visualization literacy
journalyear: 2026copyright: rightsretainedconference: Proceedings of the ACM Conference on Human Factors in Computing Systems; April 13–April 17, 2026; Barcelona, Spainbooktitle: Proceedings of the ACM Conference on Human Factors in Computing Systems (CHI ’26), April 13–April 17, 2026, Barcelona, Spaindoi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Visual analytics
@@
Figure 1. Multimodal interaction for dashboard onboarding using LLMs. Building on the process model for dashboard onboarding proposed by Dhanoa et al. (Dhanoa et al., 2022), we extend the model with multimodal interactions and LLM-powered features for user onboarding.
The figure shows how onboarding interactions, interface artifacts, and onboarding content relate within an AI-enhanced onboarding system. It shows users initiating actions such as hovering, clicking, typing, or speaking, which in turn manipulate the underlying dashboard content and activate artifacts like contextual help menus or radial menus. These artifacts are presented to the user as text, audio, and explanations generated by a language mode and this content updates based on additional user manipulation. The figure conveys a pattern of continuous feedback loops rather than a linear flow: user interactions trigger system responses, and those responses shape the next user actions, illustrating an integrated ecosystem of input, interpretation, and adaptive guidance.
1. Introduction
@@
Visualization dashboards have become essential tools for data exploration and decision-making across organizations (Sarikaya et al., 2019b). Modern dashboards integrate multiple linked views, complex filtering mechanisms, and sophisticated interaction techniques to support analytical workflows (Bach et al., 2023). However, this complexity creates a significant barrier: users often struggle to understand how to navigate and effectively use these dashboards without guidance (Walchshofer et al., 2023; Ceneda et al., 2017; Collins et al., 2018). Dashboard authors accordingly invest substantial time creating onboarding materials and conducting training sessions (Dhanoa et al., 2022; Stoiber et al., 2019), yet these resources quickly become outdated as dashboards evolve (Dhanoa et al., 2025). Existing automated onboarding solutions often fail to match the effectiveness of human experts who can provide contextual explanations, adapt to user questions, and demonstrate features using the actual data and visualizations present in the dashboard.
@@
We present Diana (Dashboard Interactive Assistant for Navigation and Analysis), a multimodal dashboard assistant that replicates the guidance provided by human onboarding experts. Unlike text-only chatbots or onboarding tutorials (static or dynamic (Dhanoa et al., 2025)), Diana combines multiple interaction modalities—voice input and output, mouse, and keyboard-based interaction—to provide situated explanations (Chundury et al., 2023) directly within the dashboard interface. The system is based on the process model for dashboard onboarding described by Dhanoa et al. (Dhanoa et al., 2022) and uses LLMs for the unique case of onboarding (Figure 1). It employs a contextual radial menu that organizes dashboard features into logical structures, supports pointing gestures and lasso selection for highlighting specific charts or data points, includes a text-based natural language interface, and integrates speech recognition and synthesis for natural conversation. Diana delivers contextual help using the actual tasks, visualizations, and datasets from each specific dashboard, enabling users to learn through direct manipulation and exploration of real content.
@@
We implemented Diana as a web-based system using a GPT-powered agent and integrated with Microsoft Power BI. The system supports any Power BI dashboard with appropriate metadata annotations. It uses OpenAI’s Realtime API (OpenAI, 2023b) for speech processing and provides real-time visual feedback through dashboard highlighting and overlays. Using the prototype, we conducted a mixed-methods user study comparing Diana against a baseline system and collected both quantitative and qualitative data from participants with novice- to expert-level dashboard experience. The study design incorporated insights from four pilot studies to refine the experimental protocol. Results demonstrate that participants with Diana had higher task accuracy as compared to the baseline. They rapidly grasped Diana’s capabilities and consistently preferred multimodal interactions over single-modality approaches for onboarding tasks. Visual highlighting combined with voice-based questions and responses emerged as the dominant interaction pattern. All experts expressed strong interest in using Diana for their professional dashboard work.
@@
The contributions of this work are threefold: (1) the concept of a multimodal dashboard onboarding assistant that combines natural language, visual highlighting, and gestural interaction to replicate human-expert guidance; (2) the implementation of Diana, demonstrating how large language models can be integrated with dashboard environments to provide contextual, situated assistance; and (3) an exploratory study comparing Diana against a standard onboarding procedure (static dashboard user guide), focusing on differences in task correctness, interaction strategies, and perceived onboarding experience.
2. Related Work
@@ 2.1. Visualization Dashboards
Visualization dashboards have become ubiquitous tools for data exploration and decision-making across diverse domains, yet their complexity often creates significant barriers for users. Few’s book on visualization dashboards (Few, 2006) outlined the many fine points and varied purposes of dashboard design in industry. Sarikaya et al. (Sarikaya et al., 2019a) provided a comprehensive survey of what the scientific visualization community means by “dashboards,” revealing their diverse applications and the varied expectations users bring to these systems. Their work highlighted the fundamental challenge: dashboards must balance powerful functionality with usability.
@@
Multiple studies have documented the struggles users face when working with dashboard systems. Tory et al. (Tory et al., 2021) examined how users attempt to find their “data voice” when working with dashboards, documenting persistent difficulties in navigation, interpretation, and effective use of interactive features. Their findings reveal that even well-designed dashboards can overwhelm users who lack sufficient training. Similarly, Walchshofer et al. (Walchshofer et al., 2023) observed the socio-technical challenges that arise when organizations transition to new dashboarding systems, noting that successful adoption depends heavily on both technical design and organizational support.
@@
The visualization community has responded to these challenges by developing design frameworks and best practices. Bach et al. (Bach et al., 2023) identified recurring dashboard design patterns, providing a vocabulary for discussing effective dashboard layouts and interaction paradigms. Lin et al. (Lin et al., 2024) developed DMiner, which provides design recommendations for dashboards based on analysis of a large dataset of existing dashboards. Their system helps authors create more effective dashboards by suggesting layouts, chart types, and design patterns that have proven successful in similar contexts, representing a shift toward data-driven dashboard design that could reduce user confusion from the outset.
@@
Despite these advances in design guidance, the challenge remains: complex dashboards with multiple coordinated views and rich interactions require significant investment of time and effort by users to master. This persistent gap between dashboard capabilities and user proficiency motivates the need for more sophisticated onboarding and assistance mechanisms that can help users navigate these powerful but complex systems.
2.2. Onboarding for Dashboards
@@
Given the popularity of visualization dashboards, visualization onboarding has recently emerged as a research area, though solutions remain fragmented across different contexts and applications. Stoiber et al. (Stoiber et al., 2019) first characterized the onboarding space for individual visualizations, establishing foundational concepts, such as when, where, and why to provide onboarding. Diana uses these concepts to provide an onboarding (where), before and during a task (when), and only when the user asks for it (why). They also identified the need for systematic approaches to help users learn visualization. Their work highlighted various onboarding methods ranging from simple tooltips to more sophisticated guidance.
@@
Building on these foundational concepts, Dhanoa et al. (Dhanoa et al., 2022) proposed a systematic process model for dashboard onboarding that identifies key stages, such as interaction with the onboarding system, manipulation of the dashboard content, and its presentaion with the help of artifacts to the users. This framework helps understand onboarding as a structured process rather than ad-hoc training. Diana uses this process to specify the interactions, content and the artifacts that are presented to the user (shown in Figure 1).
@@
Recent advances have focused on automating onboarding creation using tours. Effective guided tours balance navigational constraints and freedom to facilitate learning (Elmqvist et al., 2008). Dhanoa et al. (Dhanoa et al., 2025) introduced D-Tour, a system for semi-automatically generating interactive guided tours for visualization dashboards. Their approach reduces the manual effort required from dashboard authors while maintaining the ability to create engaging onboarding experiences. Similarly, Hoque and Sultanum (Hoque and Sultanum, 2025) developed DashGuide, which enables the creation of interactive dashboard tours with minimal author input while allowing for manual refinement when needed.
@@
The comparative evaluation of different onboarding approaches has revealed important insights about their effectiveness. Stoiber et al. (Stoiber et al., 2022) systematically compared various onboarding methods, finding that interactive approaches generally outperform static documentation, though the optimal approach depends on user characteristics and dashboard complexity.
@@
Despite these advances, most existing onboarding solutions remain primarily visual and static, relying on annotations (Elias and Bezerianos, 2012), tooltips, or guided tours that users follow passively. Chundury et al. (Chundury et al., 2023) represent a notable exception with their contextual, in-situ help features for visual data interfaces, though their approach focuses on help rather than comprehensive onboarding. This gap between the interactive nature of modern dashboards and the static nature of most onboarding solutions motivates the need for more dynamic, responsive onboarding systems that can adapt to user behavior and provide assistance through multiple interaction modalities. Current automated onboarding methods are still far inferior to the gold standard of onboarding: using a human instructor.
2.3. Multimodal Interaction with Data Visualizations
@@
The visualization community has long recognized the limitations of traditional mouse and keyboard interaction for complex data exploration tasks. Elmqvist et al. (Elmqvist et al., 2011) articulated the concept of fluid interaction for information visualization, arguing that effective visualization systems should support seamless transitions between different interaction modes and reduce cognitive barriers between user intent and system response. This foundational work established fluidity as a key principle and identified the utility of multiple modalities for designing more natural visualization interfaces. Multimodality also has the benefit of supporting accessible visualization for non-visual audiences (Chundury et al., 2022).
@@
Lee et al. (Lee et al., 2012) provided a comprehensive call to action, encouraging the visualization community to expand beyond mouse and keyboard to support more natural interactions. Their survey of alternative interaction modalities highlighted the potential for touch, speech, gesture, and other modalities to make visualization more accessible and expressive. This work catalyzed subsequent research into multimodal visualization interfaces.
@@
Early explorations of touch-based visualization interaction demonstrated the potential of direct manipulation using modalities beyond the mouse and keyboard. Baur et al. (Baur et al., 2012) introduced TouchWave, showing how kinetic multi-touch gestures could enable fluid manipulation of hierarchical stacked graphs. Their work revealed how touch affordances could support more intuitive interactions with complex visualization structures. Building on these foundations, Nielsen et al. (Nielsen et al., 2016) studied the use of touch and pen scribbling to support multidimensional visualization queries. Badam et al. (Badam et al., 2017) analyzed the affordances of different input modalities—touch, speech, proxemics, gestures, gaze, and wearables—for visual data exploration, providing guidelines for combining modalities to leverage their complementary strengths. Thompson et al. (Thompson et al., 2018) developed Tangraphe, which employed single-hand, multi-touch gestures for network visualization exploration, demonstrating how naturalistic gestures could replace conventional pointer-based interactions.
@@
The integration of speech with visualization interaction has proven particularly promising for complex analytical tasks. Srinivasan and Stasko (Srinivasan and Stasko, 2017) examined what users naturally want to express through natural language when analyzing data, providing insights into the linguistic patterns that visualization systems should support. They later developed Orko (Srinivasan and Stasko, 2018), one of the first systems to combine voice and direct manipulation for network analysis, showing how speech could complement visual interaction for complex exploration tasks. These systems demonstrated that users could express analytical intents through conversational interfaces, though challenges remained in handling ambiguous queries and maintaining context across extended analysis sessions. Building on these foundations, Srinivasan and Setlur (Srinivasan and Setlur, 2023) developed BOLT, a natural language interface specifically designed for dashboard authoring, showing how conversational interaction could extend beyond data exploration to visualization creation itself.
@@
Empirical studies have provided crucial insights into user preferences and performance with different modality combinations. Saktheeswaran et al. (Saktheeswaran et al., 2020) found that participants preferred combining multimodal over unimodal input when exploring network visualizations on large vertical displays, with touch and speech proving to be a particularly effective combination. Their findings challenged assumptions that users would prefer single-modality interaction for simplicity. In large display contexts, León et al. (Molina León et al., 2024) found that participants preferred speech commands for 10 of 15 exploration tasks over touch, pen, and mid-air gestures, suggesting that speech excels particularly in scenarios where physical interaction is challenging.
@@
Recent work has focused on designing consistent multimodal experiences across different devices. Jakobsen et al. (Jakobsen et al., 2013) explored how user position and movement relative to large displays could inform interaction design, showing how proxemic cues could enhance information visualization experiences through implicit input; Badam et al. (Badam et al., 2016) extended this to multiple users. Srinivasan et al. (Srinivasan et al., 2020) developed InChorus, which provides consistent multimodal interactions across tablet devices, demonstrating how design principles can maintain usability while adapting to different form factors.
2.4. Automatic Generation of Dashboards